OpenStudy (anonymous):

find deceleration: given 9.8 as velosity ,time= 1.7s equation vf=vi+at

OpenStudy (anonymous):

please help i need someone to double check my answer i reached to -5.7647

ganeshie8 (ganeshie8):

looks correct to me !

ganeshie8 (ganeshie8):

wats ur question ?

OpenStudy (anonymous):

they wanted me to use that answer to find yards gain

OpenStudy (anonymous):

but i received a really big answer of 44,000 plugging all the information into x=xo+vot+1/2at^2 which threw me off

ganeshie8 (ganeshie8):

xo = 0 rihgt ?

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

x = 9.8(1.7) + 1/2( -5.7647)(1.7)^2 = ?

ganeshie8 (ganeshie8):

check ur calculation, that should come ~ 8

OpenStudy (anonymous):

thank you i think i forgot the negative sign

ganeshie8 (ganeshie8):

ohhk np :)

ganeshie8 (ganeshie8):

are you doing MIT classical mechanics course by any chance ? :)

OpenStudy (anonymous):

no im doing physis. but if you need help ill try to get the answer for

ganeshie8 (ganeshie8):

sure i may need help... the course just started... il tag u when get stuck wid somthing. thank you :)

OpenStudy (anonymous):

ok anytime and thank you for your help

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

bye